Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 269
Entry time: 18 Oct 2021, 10:14
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 18.10.2021 |
|
|
General Announcement:
-
CTRL:
- Nothing to report.
DA:
- On the mention of offline calibration being slow: DA reminds to call the DOC as soon as the issue occurs. It is hard to find root causes at a later time.
EEE:
- Issue with SCS Scannery: solved during the weekend and follow up further;
- Coax cable installation in FXE to start from this Tuesday 19.10.
VAC:
- Nothing to report.
XPD:
- No issues to report; This week XPD is involved at Hard X-rays activities with 24keV in SA1 and >30keV in SA2.
DET:
- Nothing to report.
ITDM:
- Nothing to report.
XRO:
- Attempt to measure the beam position in SA2 this morning was missed as machine started to change parameters.
LAS:
- Not present.
UNSYS:
- Nothing to report.
SPB:
- Observed beam drift of up to 2mm on the 17.10; Theo M. confirms that horizontal direction moved up to 7mm at the XGM.
FXE:
- Nothing to add.
MID:
- All fine at MID; successful user beam time; Issues with Laser being helped during the weekend.
HED:
- Agipd-mini half back to HED; checked with DAQ; tests during this week.
- Also calibration for the single shot operation this week (Single shot with Agipd; picking only trains that have signal).
- Weekend to get 30keV beam: Zuzana K. is the main contact for communication and planning.
SCS:
- Scanner-y issues: it does not run stable (not allowing to be used in automated motion); Beamtime successful.
SQS:
- Nothing to report. |